# Nuclear level density of 128Te from $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering and complementary photonuclear data

### v1

@@ Abstract
We have extracted the nuclear level density of 128Te from a $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering experiment using the large-volume LaBr3:Ce and CeBr3 detectors from ELI-NP at the 9 MV Tandem facilities at IFIN-HH. The decay data were normalised using photonuclear data, resulting in nuclear level densities without intrinsic model dependencies from the constant temperature or Fermi gas models. The measured nuclear level density follows closely between the expectations from these two models, but we observe a clear divergence from a microscopic model based on the Skyrme force.
  • January 16, 2025
@@ 1 Introduction
Statistical nuclear properties, like $\gamma$-ray strength functions and nuclear level densities, are important measurable observables determining reaction rates for applications like nuclear reactor technologies and criticality, processing of spent nuclear fuel, and nuclear astrophysics. These observables serve as input to the statistical model that describes the atomic nucleus at energies and temperatures where discrete states start to overlap due to their natural widths and can not be treated as separate energy states anymore. For a recent overview of the current experimental and theoretical status of NLDs and $\gamma$SF, see Reference [1]. Several different traditional methods exist for NLD measurements. These include the well-established Oslo method [2, 3, 4, 5], evaporation spectra of protons [6] or neutrons [7], or high-energy $(\mathrm{p},\mathrm{p}^{\prime})$ scattering with magnetic spectrometers [8].
@@ 2 Experiment
The experiment was conducted at the IFIN-HH 9 MV Tandem facilities in Măgurele, Romania, in the spring of 2024. The experimental setup consisted of using a combination of the ELIGANT Gamma Neutron (ELIGANT-GN) large-volume LaBr3:Ce, and CeBr3 detectors [17, 18] mounted in the mechanical frame and inside the bismuth germanate (BGO) shields of the ROSPHERE [19], a configuration that has been previously reported in detail in Reference [20]. The total angular coverage of the scintillators in this configuration was 11.95% of the solid angle. In addition to the $\gamma$-ray detectors, a charged-particle detector array consisting of silicon-strip detectors of the type Micron S7 with thicknesses of 65 $\mu$m and 1000 $\mu$m, respectively, was placed in a $\Delta E-E$ configuration in the backward direction. These detectors were placed at a distance of 28 mm from the target for the thin detector and 44 mm from the target for the thick detector, providing an angular coverage of 122-136 in the overlap region relative to the target.
@@
The target was 1.7 mg/cm2 evaporated on a 30 $\mu$g/cm2 backing of diamond-like carbon. The isotopic purity of the target was 98.2(2)128Te with the largest impurity contribution from 130Te (1.25%), 126Te (0.37%), and 125Te (0.11%) while other isotopes of Te contributed with $<0.1$% each. On this target, a proton beam with an energy of 14 MeV and a typical beam current of $1.5$ nA impinged, and the data were collected for 70 h. The limiting factor in these measurements was the count rate in the BGO detectors. Thus, they were disabled in the data acquisition (DAQ). The DAQ system was fully digital with the LaBr3:Ce and CeBr3 detectors read out using CAEN V1730 digitisers running Digital Pulse Processing for Charge Integration and Pulse Shape Discrimination (DPP-PSD) firmware, and the silicon detectors as well as the HPGe detectors read out by CAEN V1725 digitisers running Digital Pulse Processing for Pulse Height Analysis (DPP-PHA) firmware.
@@ 3 Analysis
Figure 2: Experimental (top left) and best fit (top right) of the probability matrix. The bottom panels show two projections of the best fit (red line) compared with the experimental data (black circles) in the energy ranges 3.61-3.71 MeV and 5.31-5.41 MeV. The error bars include both statistical uncertainties and systematic uncertainties evaluated for the reduced $\chi^{2}/\mathrm{NDF}=1$.
@@
where $L$ is the multipolarity of the $\gamma$ ray and we assume $L=1$, corresponding to complete dipole domination, and $X$ correspond to either electric (E) or magnetic (M) transitions. The result of this fit is shown in Figure 2, together with two projections at selected energies to highlight the agreement between the fit and the experimental data. At this point, we evaluate the systematic uncertainties that can originate from the data processing as well as assumptions required for the method to work, from the reduced $\chi^{2}$ by taking the number of degrees of freedom (NDF) into account such that $\chi^{2}/\mathrm{NDF}=1$. The methodological systematic uncertainties could, for example, include the validity of the Brink-Axel hypothesis that the $\gamma$SFs are independent of $E_{\mathrm{x}}$ and, thus, that a unique $\gamma$SF will fit all the data. Here, we include these uncertainties and the statistical uncertainties in the analysis. We note that around $E_{\gamma}\approx 1.7$ MeV, there is an oversubtraction originating from the first-generation procedure, likely originating from a small 27Al background. This over-subtraction will propagate into the $\gamma$SF evaluation as an artificial dip. However, as the NLD contains the sum over all possible energies, it will not introduce any artificial features.
@@
By extracting $P(E_{\gamma},E_{x})$ from the measured results, we have obtained one pair of solutions for $\rho(E_{x})$ and $\mathcal{T}(E_{\gamma})$. By normalising the $\mathcal{T}(E_{\gamma})$ solution to the $(\gamma,\gamma^{\prime}\gamma^{\prime\prime})$ data [39], we can extract the parameters $B$, corresponding to the absolute values, and $\alpha$, corresponding to the slope, from equation (4), shown in Figure 3. With the slope known, we can fix the absolute value of the nuclear level densities, parameter $A$, from the complete spectroscopy of known states in the energy range 2-3 MeV, also shown in Figure 3 and listed in Table 1, under the assumption that the $\gamma$SF from the $(\gamma,\gamma^{\prime}\gamma^{\prime\prime})$ data, mainly populating a very sharp spin ($J$) and parity ($P$) distribution with $J^{P}\approx 1^{\pm}$, is close enough to the $\gamma$SF from higher spin states populated in inelastic $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering.
@@
-0.2875 0.069(22) 1.3125 0.204(24) 2.9125 1.043(10) 4.5125 1.61(28)
-0.1875 0.11(4) 1.4125 0.336(4) 3.0125 1.30(14) 4.6125 1.91(34)
-0.0875 0.21(11) 1.5125 0.471(7) 3.1125 1.49(16) 4.7125 2.16(4)
0.0125 0.24(10) 1.6125 0.454(5) 3.2125 1.53(16) 4.8125 2.17(31)
0.1125 0.138(27) 1.7125 0.471(4) 3.3125 1.83(20) 4.9125 2.62(4)
0.2125 0.106(13) 1.8125 0.535(4) 3.4125 2.12(25) 5.0125 2.64(34)
0.3125 0.0941(8) 1.9125 0.98(8) 3.5125 2.74(4) 5.1125 2.68(30)
0.4125 0.18(20) 2.0125 1.72(15) 3.6125 3.33(5) 5.2125 3.07(33)
0.5125 0.43(9) 2.1125 2.42(19) 3.7125 3.99(6) 5.3125 3.69(4)
0.6125 1.52(8) 2.2125 3.22(22) 3.8125 4.75(8) 5.4125 3.75(35)
0.7125 2.0(11) 2.3125 3.80(25) 3.9125 5.36(8) 5.5125 4.15(4)
0.8125 1.63(6) 2.4125 4.65(25) 4.0125 6.7(11) 5.6125 5.62(5)
0.9125 0.84(13) 2.5125 5.45(29) 4.1125 8.3(14) 5.7125 5.22(5)
1.0125 0.48(4) 2.6125 6.63(4) 4.2125 10.4(19) 5.8125 4.21(29)
1.1125 0.381(20) 2.7125 7.31(6) 4.3125 10.9(18)
1.2125 0.713(5) 2.8125 8.85(7) 4.4125 15.1(30)
@@ 4 Discussion
The overall agreement between the shape of the $\gamma$SF from the $(\gamma,\gamma^{\prime}\gamma^{\prime\prime})$ data [39] is shown in Figure 3. Note that since the $(\gamma,\gamma^{\prime}\gamma^{\prime\prime})$ data were used for normalisation of the absolute magnitude and the total slope of the $\gamma$SF, only the relative shapes can be used to evaluate the consistency. For comparison, we also include the following microscopic calculations and parametrisations of the $\gamma$SFs as implemented in the TALYS [41, 42] code: the temperature-dependent Hartree-Fock-Bogolyubov (HFB) by Goriely [43], the HFB and Quasiparticle Random Phase Approximation (QRPA) calculations with the Gogny D1M interaction [44], and the simplified version of the modified Lorentzian (SMLO) function from the recommendations of the International Atomic Energy Agency (IAEA) Coordinated Research Project (CRP) on $\gamma$SFs [45, 46].
On the NLD side, the procedure described in Section 3 should provide us with the NLD directly from experimental data, without the need to invoke model-dependent extrapolations to the neutron separation energy, typically via the constant temperature or the Fermi gas model. Thus, comparing the experimental data presented here to these two common extrapolations would be interesting. The two models are simple phenomenological two-parameter models that are based on the observation of empirical nuclear properties. Here, we will use the formulation of these models presented in Reference [47].
@@
This expression also consists of two main parameters, the level density parameter $a$ and the Fermi back-shift parameter $E_{1}$. In addition, the spin cut-off parameter, $\sigma$, is included explicitly in the expression (6). In fact, in both equations (5) and (6), the spin distribution has been considered independent from $E_{x}$ and separated into the function $f(J)$,
$f(J)=\frac{2J+1}{2\sigma^{2}}\exp\left[-\frac{\left(J+1/2\right)^{2}}{2\sigma^{2}}\right].$ (7)
Note that this energy independence of $f(J)$ is an approximation, and in an actual nucleus, there indeed is a likely energy dependence of $\sigma$ and, therefore, also on $f(J)$.
To compare the experimental data with the phenomenological descriptions in equations (5-7), we employed a selection of typical parametrisations of the data from Gilbert and Cameron [48], and von Egidy and Bucurescu [50]. The extraction of parameters for equations (5-7) was performed using the ROBIN code version 1.92 [51]. The obtained parameters are listed in Table 2 and the corresponding NLDs are shown in Figure 3.
@@
In addition to the phenomenological models, the calculated results from the Reference Input Parameter Library (RIPL) calculated from the Skyrme force, as implemented in TALYS 1.95, are also shown [52].

### v2

@@ Abstract
We have extracted the nuclear level density of 128Te from a $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering experiment using the large-volume LaBr3:Ce and CeBr3 detectors from ELI-NP at the 9 MV Tandem facilities at IFIN-HH. The decay data were normalised using photonuclear data, resulting in nuclear level densities without intrinsic model dependencies from the constant temperature or Fermi gas models. The deduced nuclear level density follows in between the expectations from these two models, but we observe a clear divergence from a microscopic model based on the Skyrme force.
  • April 14, 2025
@@ 1 Introduction
Statistical nuclear properties, like $\gamma$-ray strength functions and nuclear level densities, are important measurable observables determining reaction rates for applications like nuclear reactor technologies and criticality, processing of spent nuclear fuel, and nuclear astrophysics. These observables serve as input to the statistical model that describes the atomic nucleus at energies and temperatures where discrete states start to overlap due to their natural widths and can not be treated as separate energy states anymore. For a recent overview of the current experimental and theoretical status of NLDs and $\gamma$SF, see Reference [1]. Several different traditional methods exist for NLD measurements. These include well-established approaches, such as the Oslo method [2, 3, 4, 5], evaporation spectra of protons [6] or neutrons [7], or high-energy $(\mathrm{p},\mathrm{p}^{\prime})$ scattering with magnetic spectrometers [8].
@@ 2 Experiment
The experiment was conducted at the IFIN-HH 9 MV Tandem facilities in Măgurele, Romania, in spring 2024. The experimental setup consisted of using a combination of the ELIGANT Gamma Neutron (ELIGANT-GN) large-volume LaBr3:Ce, and CeBr3 detectors [19, 20] mounted in the mechanical frame and inside the bismuth germanate (BGO) shields of the ROSPHERE [21], a configuration that has been previously reported in detail in Reference [22]. The total angular coverage of the scintillators in this configuration was 11.95% of the solid angle. In addition to the $\gamma$-ray detectors, a charged-particle detector array consisting of silicon-strip detectors of the type Micron S7 with thicknesses of 65 $\mu$m and 1000 $\mu$m, respectively, was placed in a $\Delta E-E$ configuration in the backward direction. These detectors were placed at a distance of 28 mm from the target for the thin detector and 44 mm from the target for the thick detector, providing an angular coverage of 122-136 in the overlap region relative to the target.
@@
The target was 1.7 mg/cm2 and evaporated on a 30 $\mu$g/cm2 backing of diamond-like carbon. The isotopic purity of the target was 98.2(2)128Te with the largest impurity contribution from 130Te (1.25%), 126Te (0.37%), and 125Te (0.11%) while other isotopes of Te contributed with $<0.1$% each. On this target, a proton beam with an energy of 14 MeV and a typical beam current of $1.5$ nA impinged, and the data were collected for 70 h. The limiting factor in these measurements was the count rate in the BGO detectors. Thus, they were disabled in the data acquisition (DAQ). The DAQ system was fully digital with the LaBr3:Ce and CeBr3 detectors read out using CAEN V1730 digitisers running Digital Pulse Processing for Charge Integration and Pulse Shape Discrimination (DPP-PSD) firmware, and the silicon detectors as well as the HPGe detectors read out by CAEN V1725 digitisers running Digital Pulse Processing for Pulse Height Analysis (DPP-PHA) firmware.
@@ 3 Analysis
Figure 2: Experimental (top left) and best fit (top right) of the probability matrix. The bottom panels show two projections of the best fit (red line) compared with the experimental data (black circles) in the energy ranges 3.61-3.71 MeV and 5.31-5.41 MeV. For comparison, the best fit of the data without the aluminium contamination subtracted (blue dotted line) is also shown. The error bars include both statistical uncertainties and systematic uncertainties evaluated for the reduced $\chi^{2}/\mathrm{NDF}=1$.
@@
where $L$ is the multipolarity of the $\gamma$ ray and we assume $L=1$, corresponding to complete dipole domination, and $X$ correspond to either electric (E) or magnetic (M) transitions. While the total $\gamma$SF contains the sum of all $f_{XL}$, we can not explicitly separate the different multipolarity components in this data. However, measurements and theory generally support the assumption of complete dipole dominance. For example, using high-resolution forward angle inelastic proton scattering on 208Pb at the Research Center for Nuclear Physics (RCNP) in Japan, [41] showed a $\sim$ 10% contribution to the $\gamma$SF from M1 around $S_{\mathrm{n}}$, which will not influence the evaluation in Equation (2), and about 1% contribution from E2 in the 10-13 MeV region, much higher in energy than the investigation presented here. The result of this fit is shown in Figure 2, together with two projections at selected energies to highlight the agreement between the fit and the experimental data. At this point, we evaluate the systematic uncertainties that can originate from the data processing as well as assumptions required for the method to work, from the reduced $\chi^{2}$ by taking the number of degrees of freedom (NDF) into account such that $\chi^{2}/\mathrm{NDF}=1$, giving a contribution of $\sim 27$% from the statistical and $\sim 73$% from the systematic uncertainties to the total uncertainties. The methodological systematic uncertainties could, for example, include the validity of the Brink-Axel hypothesis that the $\gamma$SFs are independent of $E_{\mathrm{x}}$ and, thus, that a unique $\gamma$SF will fit all the data. Here, we include these uncertainties and the statistical uncertainties in the analysis. We note that around $E_{\gamma}\approx 1.7$ MeV, there is an oversubtraction originating from the first-generation procedure, likely originating from a small 27Al background. This over-subtraction will propagate into the $\gamma$SF evaluation as an artificial dip. However, as the NLD contains the sum over all possible energies, it will not introduce any artificial features. As this 27Al background appears in the raw matrix and, thus, has a correct reaction-kinematic reconstruction, it must originate from the target area, most likely from the target frame. While the 1.7 MeV transition is the strongest, a couple of weaker discrete transitions have been identified in Figure 1. Due to the low level density in 27Al, no statistical transitions are expected in the energy range of interest; thus, the contamination can be subtracted. The impact of this aluminium contamination is well below the statistical uncertainties in significance, even without the subtraction of the 27Al contribution, as shown in Figure 1. The 27Al contribution most prominently appears as an additional peak with an energy of 2.7 MeV and a more significant over-subtraction at 1.7 MeV.
@@
By extracting $P(E_{\gamma},E_{x})$ from the measured results, we have obtained one pair of solutions for $\rho(E_{x})$ and $\mathcal{T}(E_{\gamma})$. The normalisation procedure for obtaining the physical solution depends on the complementary information available for each case. In the regular Oslo method, the average neutron resonance spacing, $D_{0}$, is often used to obtain the NLD at the neutron threshold, $\rho(S_{\mathrm{n}})$, to fix the slope of the NLD, followed by fixing the absolute value of the $\gamma$SF to the average neutron resonance $\gamma$-ray decay width, $\langle\Gamma_{\gamma}\rangle$. In cases when the $D_{0}$ parameter is not available, other normalisation methods have been developed. The Shape method [42], for example, have been developed to begin the normalisation procedure by fixing the shape of the $\gamma$SF through decay to discrete states and use the constrained $\gamma$SF to extract the NLD. Here, we explore another option to constrain the $\gamma$SF for evaluating the NLD. By normalising the $\mathcal{T}(E_{\gamma})$ solution to the photoabsorption cross-section [43], we can extract the parameters $B$, corresponding to the absolute values, and $\alpha$, corresponding to the slope, from equation (4), shown in Figure 3. Note that we only include the photoabsorption cross-section from elastic and inelastic photon scattering from Reference [43] and not the photoneutron data from Reference [44] in the minimisation. However, both data sets are shown in Figure 3 for reference. The reason for excluding the photoneutron data is the known issue with systematic discrepancies between old photoneutron measurements between different facilities [45], which would propagate additional uncertainties into our results. The fit to the data was done using a $\chi^{2}$ minimization on the $\gamma$SF as,
$\chi^{2}=\sum_{i}\frac{({f_{(\gamma,\gamma^{\prime}),i}}-{\tilde{f}_{i})}^{2}}{\sigma_{{f_{(\gamma,\gamma^{\prime}),i}}}^{2}+\sigma_{{\tilde{f}_{i}}}^{2}},$ (5)
where $f_{(\gamma,\gamma^{\prime})}$ is the $\gamma$SF from photon scattering data and $\tilde{f}$ is the $\gamma$SF from the combination of Equations (2) and (4). The region included for the $\chi^{2}$ minimisation here was the full energy overlap region between our data and Reference [43] giving a relative uncertainty of $\sigma_{\alpha}/\alpha=3.4$% for a reduced $\chi^{2}/\mathrm{NDF}=2.65$ when corrected for the number of degrees of freedom (NDF). With the slope known, we can fix the absolute value of the NLD, parameter $A$, from the complete spectroscopy of known states in the energy range 2-3 MeV [46]. This range was chosen to have a lower energy limit above where several states start contributing to the NLD, but below the energy limit where the experimental information of the number of excited states is incomplete. The latter was chosen according to the recommendations from Reference Input Parameter Library (RIPL)-3 [47], where the recommendation of the number of levels in a complete level scheme is 110, corresponding to 3.345 MeV. The typical uncertainty in the NLD at $E_{x}>3$ MeV before normalization was $\sigma_{\rho(E_{x})}/\rho(E_{x})\approx 13$%. The contribution to the systematic uncertainty from the fitting of $\alpha$ is negligible close to $E_{x}=3$ MeV while increasing linearly for the extrapolation along $E_{x}$ for a total uncertainty of $\sigma_{\rho(E_{x})}/\rho(E_{x})\approx 50$% at the highest energy, $E_{x}=5.8$ MeV. The resulting NLD is shown in Figure 3 and listed in Table 1, under the assumption that for at the $\gamma$SF is independent of the initial spin, and only depends on the $\gamma$-ray energy and electromagnetic multipolarity in the statistical region, such that the $\gamma$SF from the photoabsorption cross-section, mainly populating a very sharp spin ($J$) and parity ($P$) distribution with $J^{P}\approx 1^{\pm}$, is equivalent to the $\gamma$SF from higher spin states populated in inelastic $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ scattering.
@@
-0.2875 0.09(4) 1.3125 0.238(35) 2.9125 1.07(6) 4.5125 1.44(35)
-0.1875 0.15(5) 1.4125 0.39(5) 3.0125 1.32(9) 4.6125 1.7(4)
-0.0875 0.27(9) 1.5125 0.54(7) 3.1125 1.50(12) 4.7125 1.9(5)
0.0125 0.31(10) 1.6125 0.52(6) 3.2125 1.53(14) 4.8125 1.9(5)
0.1125 0.18(5) 1.7125 0.53(5) 3.3125 1.80(19) 4.9125 2.3(7)
0.2125 0.14(4) 1.8125 0.60(5) 3.4125 2.08(24) 5.0125 2.3(7)
0.3125 0.119(33) 1.9125 1.09(8) 3.5125 2.67(35) 5.1125 2.3(7)
0.4125 0.23(6) 2.0125 1.89(11) 3.6125 3.2(5) 5.2125 2.6(9)
0.5125 0.54(13) 2.1125 2.64(12) 3.7125 3.8(6) 5.3125 3.1(11)
0.6125 1.9(4) 2.2125 3.49(12) 3.8125 4.5(8) 5.4125 3.1(11)
0.7125 2.5(5) 2.3125 4.08(11) 3.9125 5.0(9) 5.5125 3.4(13)
0.8125 2.0(4) 2.4125 4.96(9) 4.0125 6.3(12) 5.6125 4.6(17)
0.9125 1.02(20) 2.5125 5.76(10) 4.1125 7.7(15) 5.7125 4.2(18)
1.0125 0.57(12) 2.6125 6.95(17) 4.2125 9.6(20) 5.8125 3.4(16)
1.1125 0.45(9) 2.7125 7.59(30) 4.3125 9.9(22)
1.2125 0.84(15) 2.8125 9.1(4) 4.4125 13.6(32)
The general features of the experimental spin distribution can be evaluated from the HPGe detectors. For this purpose, an example of a HPGe spectrum with an excitation energy gate between $9.0127.762<E_{x}<$ 9.012 MeV is shown in Figure 4, showing the characteristic peaks associated with the $J^{\pi}=6^{+}\to 4^{+}$, $J^{\pi}=4^{+}\to 2^{+}$, and $J^{\pi}=2^{+}\to 0^{+}$ transitions.
Figure 4: (Left) HPGe spectrum showing the low-lying transitions in the yrast band, highlighted in red, from cascade $\gamma$ rays, and identified non-yrast transitions, with an excitation energy gate around the neutron separation threshold. (Right) Intrinsic nuclear spin distribution at the neutron threshold from three models shown with solid lines: the Rigid Moment of Inertia and the Fermi Gas with and without empirical odd-even staggering. The estimate of the experimental spin distribution from the HPGe detectors, consisting of the intrinsic spin distribution folded by the relative population cross-section, is shown as red points for comparison.
We can estimate the spin distribution $S(J)$ at spin $J$, from the efficiency corrected intensity of the $\gamma$-ray transitions, $I_{\gamma}$ populating and depopulating states with spin $J$ [48],
$S(J)=I_{\gamma}(J\to J-2)-I_{\gamma}(J+2\to J).$ (6)
While this method gives stronger constraints for well-deformed nuclei with a clear population of high-spin values, it can still add some constraints to the spin range populated in this work. The subtracted and normalised spin distributions are shown in Figure 4 together with three theoretical models based on
$S(J)=\frac{2J+1}{2\sigma^{2}}\exp\left[-\frac{\left(J+1/2\right)^{2}}{2\sigma^{2}}\right].$ (7)
Figure 4 shows the distribution for two values of $\sigma$, the Rigid Moment of Inertia (RMI) with $\sigma_{\mathrm{RMI}}=6.263$ and the Fermi Gas model from von Egidy and Bucurescu [49] with $\sigma_{\mathrm{E\&B(FG)}}=4.414$. The third model is the Fermi Gas model from von Egidy and Bucurescu with an additional staggering term added [50] as $S^{\prime}(J)=S(J)(1+x)$ where $x=1.02$ for $J=0$, $x=0.227$ for even $J$ and $J=-0.227$ for odd $J$. It is important here to point out that these models provide us with the predicted intrinsic spin distribution of the statistical states, while the experimental data only give an estimate of the experimental spin distribution, meaning the intrinsic spin distribution is folded by the spin dependency of the population cross-section. The most extreme case in this work is the photonuclear data where almost the entire cross-section is in $J=1$, but also for the $(\mathrm{p},\mathrm{p}^{\prime}\gamma)$ data, the spin-dependency of the population cross-section is expected to narrow down the experimental spin distribution [48].
@@ 4 Discussion
The overall agreement between the shape of the $\gamma$SF from the photoabsorption cross-section [43] is shown in Figure 3. Note that since the photoabsorption cross-section was used to normalise the absolute magnitude and the total slope of the $\gamma$SF, only the relative shapes can be used to evaluate the consistency. For comparison, we also include the following microscopic calculations and parametrisations of the $\gamma$SFs as implemented in the TALYS [51, 52] code: the temperature-dependent Hartree-Fock-Bogolyubov (HFB) by Goriely [53], the HFB and Quasiparticle Random Phase Approximation (QRPA) calculations with the Gogny D1M interaction [54], and the simplified version of the modified Lorentzian (SMLO) function from the recommendations of the International Atomic Energy Agency (IAEA) Coordinated Research Project (CRP) on $\gamma$SFs [45, 55].
On the NLD side, the procedure described in Section 3 will provide us with the NLD directly from experimental data without the need to invoke model-dependent extrapolations to the neutron separation energy, typically via the constant temperature or the Fermi gas model. Thus, comparing the experimental data presented here to these two common extrapolations would be interesting. The two models are simple phenomenological two-parameter models that are based on the observation of empirical nuclear properties. Here, we will use the formulation of these models presented in Reference [56].
@@
This expression also consists of two main parameters, the level density parameter $a$ and the Fermi back-shift parameter $E_{1}$. In addition, the spin cut-off parameter, $\sigma$, is included explicitly in the expression (9). In fact, in both equations (8) and (9), the spin distribution has been considered independent from $E_{x}$ and separated into the function $S(J)$ from Equation (7). Note that this energy independence of $S(J)$ is an approximation. In an actual nucleus, there is a likely energy dependence of $\sigma$ and, therefore, also on $S(J)$.
To compare the experimental data with the phenomenological descriptions in equations (8-7), we employed a selection of typical parametrisations of the data from Gilbert and Cameron [57], and von Egidy and Bucurescu [49]. The extraction of parameters for equations (8-7) was performed using the ROBIN code version 1.92 [59]. The obtained parameters are listed in Table 2, and the corresponding NLDs are shown in Figure 3.
@@
In addition to the phenomenological models, the calculated results from the RIPL calculated from the Skyrme force, as implemented in TALYS 1.95, are also shown [60].